Compare and contrast the trajectories of a diver leaving the 10m platform with
A running take-off with a horizontal velocity of 3.0 ms-1
1
Expert's answer
2021-03-02T18:04:31-0500

1)


"x=3t,y=10+2t-0.5(9.8)t^2"


2)


"x=3\cos{60}t,y=10+3\sin{60}t-0.5(9.8)t^2"